OpenStudy (anonymous):

I need someone to walk me through the sqrt of 10x times the sqrt of 2x (i need to simplify)

OpenStudy (campbell_st):

ok... so the problem is \[\sqrt{10x} \times \sqrt{2x}\]

OpenStudy (campbell_st):

there is a law for radicals that says \[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\] you need to use this law